A population consists of 1, 5, 6, 8, 12, 7, and 11. Suppose a sample of size 3. Find the 

mean and variance.


1
Expert's answer
2022-05-12T04:18:55-0400

Mean of population 

"\mu=\dfrac{1+5+6+8+12+7+11}{7}=\dfrac{50}{7}""\approx7.142857"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}"


"=\dfrac{1}{7}((1-\dfrac{50}{7})^2+(5-\dfrac{50}{7})^2+(6-\dfrac{50}{7})^2"

"+(8-\dfrac{50}{7})^2+(12-\dfrac{50}{7})^2+(7-\dfrac{50}{7})^2"

"+(11-\dfrac{50}{7})^2)=\dfrac{580}{49}\approx11.836735"

Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\dfrac{50}{7}=\mu"


Variance of sampling distribution (without replacement)


"Var(\bar{X})=\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"=\dfrac{580}{49(7)}(\dfrac{7-2}{7-1})=\dfrac{1450}{1029}"